OpenStudy (anonymous):

What is the length of the leg in a right triangle whose hypotenuse is 13 and other leg is 12. A. 3 B. 5 C. 4 D. 2

OpenStudy (anonymous):

5!

OpenStudy (anonymous):

B= 5 do you need an expo

OpenStudy (anonymous):

here is the solution in brief according to pythagorus theorem 13(square)=12(square)+ x(square) 169=144+x(square) 169-144=x(square) 25=x(square) 5=x

OpenStudy (anonymous):

the last step is square root

OpenStudy (anonymous):

thanks guys!